A poll of 200 registered voters was taken before the election for mayor of Springdale. All 200 voters indicated which 1 of the 4 candidates they would vote for. The results of the poll are given in the table below.
Candidate Number of voters
Blackcloud 50
Lue 80
Gomez 40
Whitney 30
1) What percent of the voters polled chose Whitney in the poll?
A. 15% B. 20% C. 25% D. 30% E. 40%
2) If the poll is indicative of how the 10,000 registered voters of Springdale will actually vote in the election, which of the following is the best estimate of the number of votes Lue will receive in the election?
a) 1,500 b) 2,500 c) 4,000 d) 5,000 e) 8,000
3)If the information in the table were converted into a circle graph (pie chart), then the central angle of the sector for Gomez would measure how many degrees?
A. 54° B. 72° C. 90° D. 108° E. 144°

Answer by richwmiller(17219)

1) 30*100/200=15%
2) 80/200=40/100=.4
(40/100)*10000=40*100=4000
3)40/200=.2
.2*360=72 degrees